OpenStudy (anonymous):

Please help :) Determine the slope of the line. x y 2 1 4 6 6 11 8 16 10 21 12 26 14 31 Using the slope, write the equation o the linear function. Write it in point-slope.

OpenStudy (anonymous):

ok so do u know the formula?

OpenStudy (anonymous):

for slope that is

OpenStudy (anonymous):

yeah i do! but i tried testing it out with different ordered pairs and am getting different slopes.

OpenStudy (anonymous):

hmm lets see so for the first two pairs heres the slope:

OpenStudy (anonymous):

wait, i think i did something wrong oops let me double check

OpenStudy (anonymous):

oops haha i got 5/2

OpenStudy (anonymous):

ok good :3

OpenStudy (anonymous):

Yes it is 5 /2

OpenStudy (anonymous):

what do i do for point slope? i know the equation, but do i just use any order pair?

OpenStudy (anonymous):

ordered pair*

OpenStudy (anonymous):

formula=y2-y1\[y2-y1/x2-x1\]

OpenStudy (anonymous):

yes u can use any order paired

OpenStudy (anonymous):

so would it just be y-1=5/2(x-2) or y-1=5/2x-5

OpenStudy (anonymous):

is that how i write the equation of the linear function??

OpenStudy (anonymous):

isn't that slope intercept

OpenStudy (anonymous):

y – y1 = m(x – x1) srry lol

OpenStudy (anonymous):

haha its oK! so is my answer right? is that all i need for the equation?

OpenStudy (anonymous):

ya its right

OpenStudy (anonymous):

okay thanks!!

OpenStudy (anonymous):

i knew u could do it lol